bn converges 18. Let (an)n=1 and (bn)n=1 be sequences in R. Show that if and lan – an+1 < oo, then anbr converges.
Answer #1

let Sn = 91 + a₂ + antan let an = a bit az byt - tambn Then, to = sobi + (S2-51) b2 + (S3-S₂) b3 +---+ (Sn-Snap bn. = sebi-by